Solve the following differential equations:
i) dy/dx = cot(y + x) −1
1
Expert's answer
2019-04-30T12:22:53-0400

Solution.  We write the equation as


"\frac {dy} {dx} =\frac {1} {tan(x+y)} - 1"

"\frac {dy} {dx} =\frac {1-tan(x+y} {tan(x+y)}"

"tan(x+y)dy=(1-tan(x+y))dx"

"(tan(x+y)-1)dx+tan(x+y)dy=0"

Consider the equation as


"M(x,y)dx+N(x,y)dy=0."

Therefore


"M(x,y)=tan(x+y)-1"

"N(x,y)=tan(x+y)"

"\frac {\partial M} {\partial y}= \frac {1} {cos^2 (x+y)}"

"\frac {\partial N} {\partial x}= \frac {1} {cos^2 (x+y)}""\frac {\partial M} {\partial y} = \frac {\partial N} {\partial x}"

This equation is an equation in full differentials. Let U(x,y) is solution of the equation. Therefore


"\frac {\partial U} {\partial x}= tan(x+y)-1= \frac {sin(x+y} {cos(x+y)}-1"

Integrating x we get


"U(x,y)=-\ln cos(x+y)-x+C(y)"


where C(y) is function of y. Hence

"\frac {\partial U} {\partial y}= tan(x+y)+C'(y)= tan(x+y)"

"C'(y)=0"

"C(y)=C"

where C is constant. Therefore get


"U(x,y)=-\ln cos(x+y)-x+C"

where C is constant.

Answer.


"U(x,y)=-\ln cos(x+y)-x+C"

where C is constant.